# The Schwurbelarchiv: a German Language Telegram dataset for the Study of Conspiracy Theories

### v2

@@ 3.0.4 Ethics and Data Protection
While the data contained in the Schwurbelarchiv was collected from publicly accessible chats, it should not be considered “fair game” for research without consideration for the privacy expectations of the users whose messages are contained in the dataset. In preparing the dataset for publication, we considered a number of arguments to weigh the potential infringement on user expectations against the merit of potential research enabled by making the data accessible, which we will detail in this section. Furthermore, we obtained IRB clearance for research with and re-publication of the data contained in the Schwurbelarchiv (vote of the IRB of Graz University of Technology from March 28, 2023).
@@
Telegram users can freely choose their username on the platform. Therefore, more than one user can choose the same username and usernames do not necessarily allow the identification of individuals outside Telegram. Nevertheless, many usernames contained in the meta information of posts resemble names of real people. Even though the content contained in the dataset is public, users of social media platforms usually do not expect their posts to be singled out and discussed in research projects Townsend and Wallace (2016). We therefore argue that any attempt at linking Telegram posts to the identities of people outside the platform would be problematic from the perspective of user expectations. Therefore, we pseudonymised all messages by replacing usernames with unique IDs via hashing. As a result, messages that come from the same username can still be linked, but any resemblance of usernames to real names is removed. We believe that this step does not hinder the exploration of discussion topics and communication dynamics while it helps to balance user expectations with the potential merit of using the Schwurbelarchiv for research.
@@
A similar consideration has to be made for the names of groups and channels in the dataset. While the majority of these names refer to places or discussion topics, a number of chat names resemble the names of real people. This is particularly frequent for channels where single individuals post content that is broadcast to the followers of the channel. For the present data publication, we chose to retain the names of chats and not pseudonymise them. This decision is driven by two main considerations: First, by altering chat names, analyses like the identification of the geographical scope of chats provided in this work would not be possible and our work would not be reproducible. Second, users that choose to create a chat under their name expect to receive a substantial amount of attention. We argue that it is ethically admissible to retain these names Townsend and Wallace (2016), as providing them in the dataset will neither significantly heighten the attention these users receive nor breach their expectations.
@@
Even though usernames are anonymised in our dataset, the content of the posts might allow for the re-identification of people based on disclosed attributes such as location, demographic characteristics or real names that are contained in message texts rather than usernames. Pseudonymisation of this information in message texts let alone multimedia content is not technically feasible. However, re-identification of individuals would require some effort to parse content and link user identities to messages. While we consider this technically feasible for individual messages, we caution against any such attempts to honour expectations of individual users.
While the anonymous collector of the Schwurbelarchiv explicitly states their intention for the usage of the dataset as a source for journalists and research on the website of the dataset [^4] https://web.archive.org/web/20241126233137/
https://schwurbelarchiv.wordpress.com/
@@
, neither the website nor the Internet Archive collection contain a license. This in principle precludes re-use and re-publication of the dataset. However, in private conversation with the authors of the present work, the anonymous collector confirms that they intended to publish their data under a CC0 license and explicitly allow for re-use and re-publication.
@@ 4.1 Descriptive Statistics
Table 1 shows descriptive statistics of the cleaned Schwurbelarchiv dataset. With the number of channels and groups being almost the same, our dataset can for example be used to investigate whether conspiracy theory related discussion arises organically through conversation in discussion groups, or whether agenda-setting by a small number of users in broadcast channels is more prevalent. We find that media files like images, videos, and voice messages make up 26.4% of all messages. While they account for 38.1% of all forwarded messages, for original messages only 16.9% are media files. This difference makes the analysis of the content of media files particularly relevant for investigating content dissemination. The distinction between forwarded messages and original messages also serves to estimate the scope of the conspiracy related German-language Telegram discourse, which we will leverage in Section “Validation and Completeness” below to gauge the completeness of the dataset.
@@
Chats 6076
Authors 357,048 96,037 402,549
Channels 3130
Groups 2946
Messages total 56,230,491
Text 36,097,935 12,612,543 48,710,478
Images 5,707,134 5,994,015 11,701,149
Video 1,066,416 1,408,752 2,475,168
Voice messages 543,077 345,803 888,880
Table 1: Descriptive statistics of the content of the Schwurbelarchiv.
@@ 4.2 Language, time zone and region
To evaluate the language distribution within the chats in the dataset, we classify all messages with the FastText package Facebook (2016). Filtering out empty, short (less than 20 characters) and URL-only messages we are left with 52% of all messages. From those, German is by far most prevalent, with 92.9% of the messages followed by English with 1.8%. We also examined the distribution of German message proportions across chats, shown in Figure 1. The vast majority of chats (88.7%) contain at least 50% of messages in German, while in 46.9% of chats it exceeds 90%.
@@
As messages on Telegram are not geo-tagged, we have no direct information about the geolocation of the authors that posted messages contained in the dataset. However, we can use other information to narrow the geographic scope of the observations. In Fig. 2 we show the posting pattern on different weekdays and hours of a day using the UTC+1 time zone (e.g., Central Europe). We observe a consistent minimum in posting frequency during the European night hours with maxima in posting frequency in the early evening hours and on Sundays. This observation indicates that the diurnal rhythm of the majority of authors in the dataset is consistent with Central Europe.
Figure 2: Posting frequency aggregated by hour of the day and day of the week in the UTC+1 time zone.
We also conducted this analysis for the chat names included in the datasets provided by Zehring and Domahidi (2023) and Mohr (2023). Table 2 compares the regional representation across these datasets, showing the total number and percentages of regionally related chats. The Schwurbelarchiv dataset has the highest proportion of regionally related chats, accounting for 25%, compared to 6% and 10% in the other datasets. The high proportion of regional chats in the Schwurbelarchiv dataset highlights its value for the study of localised communication and group dynamics, as opposed to discourse dominated by large national or transnational political actors or influencers.
Number of Regional Related Chats % of full dataset
Schwurbelarchiv 1,514 24.9 %
Mohr 2,275 9.9 %
Zehring 33 5.7 %
Table 2: Comparison of regional Telegram chats across datasets.
@@ 4.3 Validation and Completeness of the Dataset
In the light of frequent message deletions on Telegram Buehling (2024) and the absence of a structured database of Telegram content that can be used to retrieve a comprehensive dataset for a given query, it is important to assess the completeness of collected Telegram datasets. To this end, we consider completeness on three levels. On the level of messages, we assess the fraction of deleted content. On the level of forwarded messages, we assess the completeness of the dataset by analysing the presence of original messages and forwarded messages: if many forwarded messages are not present in the dataset as original messages, this is evidence that a substantial part of the discourse is not captured. Lastly, we assess completeness on the level of chats included in the dataset: there is no official record of all German-speaking Telegram chats dedicated to conspiracy-related content. We therefore compare the Schwurbelarchiv with similar datasets, specifically those by Zehring and Domahidi (2023), Mohr (2023) and Baumgartner et al. (2020) introduced earlier.
@@ 4.3.1 Deleted Messages
Assessing the completeness of the dataset on a message level involves examining the potential for missing or deleted messages, which is closely tied to the data collection approach. There is no published metadata on the time of data collection. In particular, the start of the streaming process for each chat is not known. However, the difference between bulk download and streaming has direct consequences for the expected completeness of messages in the dataset at different points in time: In the bulk download phase, it is probable that the dataset is missing messages that were deleted by authors prior to the download and that the percentage of missing messages increases further back in time from the time of the download. These deleted messages have no placeholders in the dataset and are therefore not visible. In contrast, during the continual streaming phase, messages are captured in near real-time, therefore reducing the probability of missing messages caused by user deletions. We were not able to detect the transition point where the data collection method changed from bulk to stream download for the different chats. Therefore, we cannot guarantee the completeness of the dataset on a message level, since we cannot indicate the presence or fraction of deleted messages. However, it is probable that message deletions are less likely to have occurred for the later years contained in the dataset.
As stated by Buehling (2024), researchers should consider the systematic bias and possibly skewed results in computational analysis introduced by message deletions when working with such datasets. These findings indicate a key limitation of our dataset, since the lack of metadata detailing the data collection process prevents us from determining the extent to which deleted messages may have influenced its completeness.
@@ 4.3.2 Forwarded Messages
To further assess the completeness of conspiracy-related chats in the dataset, we performed a comparative analysis of forwarded messages and messages. In the context of Telegram, a forwarded message refers to a message which has been forwarded and is marked as such in the chat. These forwarded messages often serve as proxies for content influence, reflecting the flow of information between chats. We calculated the proportion of forwarded messages which are also present in the dataset as original messages. This metric indicates the degree to which the original sources of forwarded messages are represented in the dataset. If this value were 100%, it would imply that all forwarded messages are also present in the original chat the message was posted in, achieving complete coverage of the source population. In our dataset, this proportion was calculated to be 47.9%. This finding suggests that we cover about half of the German-language conspiracy related discourse on Telegram in the observation time frame, highlighting the dataset’s limitations in capturing the full extent of the discourse.
4.3.3 Chats
@@
Lastly, we also compare the contents of the Schwurbelarchiv to the chats collected by Zehring and Domahidi (2023), Mohr (2023) and Baumgartner et al. (2020). The dataset by Zehring and Domahidi (2023), consists of 578 chats and covers the time period October 28, 2015, to January 3, 2022. Among these chats, there are two that we have excluded from our dataset due to their classification as dump chats. Out of the remaining chats, 121 are also present in our dataset, representing approximately 22% of Zehring’s chats and 2% of the chats contained in the Schwurbelarchiv. Although this overlap is smaller than expected given the similar selection criteria and seed topics, several factors may account for the discrepancy: Zehring and Domahidi (2023) concentrated their seed selection on groups with substantial societal influence, including opinion leaders, politicians, and well-known conspiracy theorists. In contrast, the Schwurbelarchiv dataset includes numerous smaller, more specialized groups often focused on local issues. For example, groups such as Corona Rebellen have multiple sub-groups tailored to different cities and regions, reflecting a tendency towards decentralization in the Schwurbelarchiv dataset. This is also reflected in the overall share of content we identified as coming from regional groups, which is much higher than for Zehring and Domahidi (2023) (see also Tab. 2).
@@
The subset of German-language chats in the dataset collected by Mohr (2023) which was used for comparison here consists of 22,968 chats. Among these, 46 have been categorised by us as dump chats and have been excluded. Another 801 of the chats in this dataset are also contained in the Schwurbelarchiv. This corresponds to 13% of the chats we have in our dataset, while only comprising 3% of chats contained in the dataset collected by Mohr (2023).
@@
The Pushshift dataset Baumgartner et al. (2020) is the most comprehensive public dataset to date, encompassing over 27,000 chats from various countries. Of these, 78 chats (1.3%) overlap with the Schwurbelarchiv. None of the excluded dump chats in the Schwurbelarchiv are present in the Pushshift dataset. This low overlap is likely due to linguistic and regional differences in the selection criteria: while the Schwurbelarchiv and datasets by Zehring and Domahidi (2023) and Mohr (2023) focus on German-speaking chats, the Pushshift dataset has a more global scope. In addition, the temporal scope of data collection between the Pushshift data set (2015-2019) and the Schwurbelarchiv (2016-2022) differs substantially.
Number of Chats Overlap with Schwurbelarchiv Overlap in %
Mohr (2023) 22,968 801 13 %
Zehring and Domahidi (2023) 578 121 2 %
Baumgartner et al. (2020) 27,801 78 1.3 %
@@
Table 3 summarizes the number of chats in each dataset and their overlap with the Schwurbelarchiv, reported both in absolute numbers and as a percentage of Schwurbelarchiv chats. Furthermore, the overlap between the Zehring and Domahidi (2023) and Mohr (2023) datasets is 369 chats, corresponding to 66% and 1.6% of the content of each of these datasets, respectively. A total of 102 chats are present in all three data sets. Additionally, 39 chats are shared between the Mohr (2023) and Pushshift datasets, while no chats overlap between the datasets by Zehring and Domahidi (2023) and Pushshift.
Figure 3: Number of messages (blue) and active authors (red) per day in the Schwurbelarchiv. Maxima in messaging activity and related real-life events were identified manually.
Figure 4: Number of messages in the time period between January 1, 2022 and the end of the observation period (blue) and number of messages posted per person per day (green).
4.4 Overall Posting Activity
Figure 3 shows the overall posting frequency in terms of messages and active chats over the full observation period. Active chats in this context are defined as chats that had at least one posted message on a given day. The number of active chats and message volume already started to increase slightly before the onset of the COVID-19 pandemic. However, the dramatic increase in active users and messages coincides with the onset of the pandemic in Europe. Peaks in message volume coincide with political events such as a large demonstration at the German parliament on August 29, 2022, and the Capitol Riot in the U.S. on January 6, 2021. Further insight into the discussion topics at these times could be gained by employing topic modelling approaches to the messages posted during these events.
In Fig. 4 we again show the overall number of messages per day, overlaid with the number of messages per author and day. After an initial period of high activity in early 2020 with around 13 messages per author and day we observe a steep decline of individual author activity in April and May 2020 that does not coincide with a decline in the overall posting traffic. This indicates that while individual authors became less active, more authors joined the discourse on Telegram, contributing to a steady increase of overall message volume. This is also consistent with our impressions from deep-reading small fractions of the content: we find that as conspiracy-related Telegram chats gain in popularity, established authors are replaced by newcomers that want to learn about conspiracy theories like QAnon, Pizzagate, or Chemtrails, but rarely establish themselves as experts. We call this behaviour Conspiracy Tourism which might be an interesting avenue for further more qualitative research. We observe a more gradual decline in author activity from June 2020 to June 2021 where authors post on average nine messages a day, followed by another steep decline in late summer 2021. At the end of our observation period, e.g. in summer 2022, the activity of authors seems to have stabilised around three messages per author and day.
4.5 Activity distribution
Figure 5: A. Distribution of messages and authors per chat. Full lines indicate the logbinned distributions. B. Cumulative fraction of messages per top chats and authors. $G$ indicates the Gini coefficient of the cumulative distribution.
A common aspect of social media platforms is high inequality, where few drivers (authors, chats) are responsible for most of the content/attention. In Fig. 5 we show this to be true in the Schwurbelarchiv as well. The probability distributions of messages and unique authors per chat (Fig. 5A) have very wide ranges, with the largest chats having on the order of $10^{4}$ authors and $10^{6}$ messages. At the same time, most chats have less than $10$ messages and authors. Due to the uncertainty around the dataset sampling scheme, we refrain from e.g. fitting power-law distributions as the sampling can heavily bias both measured exponents and model selection of the best-fitted distribution (power-law vs lognormal). Regardless, the distributions are clearly heavy-tailed.
Grouping the messages by unique authors and chats (Fig. 5B), we see that $80\%$ of all messages belong to the top $12.1\%$ and $3.1\%$ of chats and authors, respectively. This is reflected in the very high Gini indexes of $G=0.85$ (chats) and $G=0.94$ (authors). This high level of concentration is comparable with what is observed in other social media platforms Orellana-Rodriguez and Keane (2018); Zhu and Lerman (2016).

### v3

@@ 3.3 Ethics and Data Protection
While the data contained in the Schwurbelarchiv was collected from publicly accessible chats, it should not be considered “fair game” for research without consideration for the privacy expectations of the users whose messages are contained in the dataset. In preparing the dataset for publication, we considered a number of arguments to weigh the potential infringement on user expectations against the merit of potential research enabled by making the data accessible, which we will detail in this section.
@@
Telegram users can freely choose their username on the platform. Therefore, more than one user can choose the same username and usernames do not necessarily allow the identification of individuals outside Telegram. Nevertheless, many usernames contained in the meta information of posts resemble names of real people. Even though the content contained in the dataset is public, users of social media platforms usually do not expect their posts to be singled out and discussed in research projects [16]. We therefore argue that any attempt at linking Telegram posts to the identities of people outside the platform would be problematic from the perspective of user expectations. Therefore, we pseudonymized all messages by replacing usernames with unique IDs via hashing. As a result, messages that come from the same username can still be linked, but any resemblance of usernames to real names is removed. We believe that this step does not hinder the exploration of discussion topics and communication dynamics while it helps to balance user expectations with the potential merit of using the Schwurbelarchiv for research.
@@
A similar consideration has to be made for the names of groups and channels in the dataset. While the majority of these names refer to places or discussion topics, a number of chat names resemble the names of real people. This is particularly frequent for channels where single individuals post content that is broadcast to the followers of the channel. For the present data publication, we chose to retain the names of chats and not pseudonymize them. This decision is driven by two main considerations: First, by altering chat names, analyses like the identification of the geographical scope of chats provided in this work would not be possible and our work would not be reproducible. Second, users that choose to create a chat under their name expect to receive a substantial amount of attention. We argue that it is ethically admissible to retain these names [16], as providing them in the dataset will neither significantly heighten the attention these users receive nor breach their expectations.
@@
Even though usernames are anonymized in our dataset, the content of the posts might allow for the re-identification of people based on disclosed attributes such as location, demographic characteristics or real names that are contained in message texts rather than usernames. Pseudonymization of this information in message texts let alone multimedia content is not technically feasible. However, re-identification of individuals would require some effort to parse content and link user identities to messages. While we consider this technically feasible for individual messages, we caution against any such attempts to honour expectations of individual users.
@@
While the anonymous collector of the Schwurbelarchiv explicitly states their intention for the usage of the dataset as a source for journalists and research on the website of the dataset [^9] https://web.archive.org/web/20241126233137/, neither the website nor the Internet Archive collection contain a license. This in principle precludes re-use and re-publication of the dataset. However, in private conversation with the authors of the present work, the anonymous collector confirms that they intended to publish their data under a CC0 license and explicitly allow for re-use and re-publication.
Furthermore, we obtained IRB clearance for research with and re-publication of the data contained in the Schwurbelarchiv (vote of the Institutional Review Board of Graz University of Technology from March 28, 2023).
@@ 4 Results
In the following section, we will first describe the Schwurbelarchiv in more detail, giving statistics on the different content types it contains as well as linguistic and temporal markers. In addition, we provide an analysis of overall posting activity and activity distribution across authors and chats. After establishing basic characteristics of the data set, we describe our efforts to assess the completeness of the data set and the extent to which it relates to conspiracy theories. Lastly, we provide more detail on the results of the audio content transcription.
@@ 4.1 Descriptive Statistics
Table 2 shows descriptive statistics of the cleaned Schwurbelarchiv dataset. With the ratio of channels to groups in our dataset being roughly 2:1  [^10] We use HTML-parsing heuristics to determine the category of a chat, which resulted in a small number of chats (37) to be manually determined., our dataset can for example be used to investigate whether conspiracy theory related discussion arises organically through conversation in discussion groups, or whether agenda-setting by a small number of users in broadcast channels is more prevalent. We find that media files like images, videos, and voice messages make up 24.4% of all messages. While they account for 42.2% of all forwarded messages, for original messages only 16.7% are media files. This difference makes the analysis of the content of media files particularly relevant for investigating content dissemination. The distinction between forwarded messages and original messages also serves to estimate the scope of the conspiracy related German-language Telegram discourse, which we will leverage in Section “Validation and Completeness” below to gauge the completeness of the dataset.
@@
Total chats 5,807
Channels 3,486
Groups 2,321
Authors 348,812 611,164 896,339
Content 43,739,502 18,855,979 62,595,481
Text 38,021,236 15,797,326 53,818,562
Images 5,707,321 6,166,370 11,873,691
Video 1,065,146 1,431,740 2,496,886
Voice messages 542,928 354,531 897,459
External links 6,965,776 5,356,078 12,321,854
Table 2: Descriptive statistics of the content of the Schwurbelarchiv. A small number of chats (15) were of an unknown category.
@@ 4.2 Language, time zone and region
To evaluate the language distribution within the chats in the dataset, we classify all messages with the Lingua package. Filtering out empty, short (less than 5 characters, e.g. emoji replies) and URL-only messages we are left with 52.5% of all messages. From those, German is by far most prevalent, with 91.3% of the messages followed by English with 2.2%. We also examined the distribution of German message proportions across chats, shown in Figure 1A. The vast majority of chats (98.6%) contain at least 50% of messages in German, while in 85.5% of chats it exceeds 90%.
@@
As messages on Telegram are not geo-tagged, we have no direct information about the geolocation of the authors that posted messages contained in the dataset. However, we can use other information to narrow the geographic scope of the observations. In Fig. 1B we show the posting pattern on different weekdays and hours of a day using the UTC+1 time zone (e.g., Central Europe). We observe a consistent minimum in posting frequency during the European night hours with maxima in posting frequency in the early evening hours and on Sundays. This observation indicates that the diurnal rhythm of the majority of authors in the dataset is consistent with Central Europe.
Figure 1: A. Percentage of German messages in chats. B. Posting frequency aggregated by hour of the day and day of the week in the UTC+1 time zone. C. Frequency distribution of messages and authors per chat. G indicates the Gini coefficient of the cumulative distribution. D. Lorenz plot showing cumulative fraction of messages per top chats and authors.
4.3 Overall Posting Activity
Figure 2A shows the overall posting frequency in terms of messages and active chats over the full observation period. Active chats in this context are defined as chats that had at least one posted message on a given day. The number of active chats and message volume already started to increase slightly before the onset of the COVID-19 pandemic. However, the dramatic increase in active chats and messages coincides with the onset of the pandemic in Europe. Peaks in message volume coincide temporally with political events such as a large demonstration at the German parliament on August 29, 2020, and the Capitol Riot in the U.S. on January 6, 2021. However, these associations are based on temporal co-occurrence and have not been verified through topic modelling or similar content analyses. Further insight into the discussion topics at these times could be gained by employing topic modelling approaches to the messages posted during these events.
In Fig. 2B we again show the overall number of messages per day, overlaid with the number of messages per author and day. After an initial period of high activity in early 2020 with around 13 messages per author and day we observe a steep decline of individual author activity in April and May 2020 that does not coincide with a decline in the overall posting traffic. This indicates that while individual authors became less active, more authors joined the discourse on Telegram, contributing to a steady increase of overall message volume. This is also consistent with our impressions from deep-reading small fractions of the content: we find that as conspiracy-related Telegram chats gain in popularity, established authors are replaced by newcomers that want to learn about conspiracy theories like QAnon, Pizzagate, or Chemtrails, but rarely establish themselves as experts. We call this behaviour Conspiracy Tourism which might be an interesting avenue for further, more qualitative research. We observe a more gradual decline in author activity from June 2020 to June 2021 where authors post on average nine messages a day, followed by another steep decline in late summer 2021. At the end of our observation period, e.g. in summer 2022, the activity of authors seems to have stabilized around three messages per author and day.
Figure 2: A. Number of messages (blue) and active chats (red) per day in the Schwurbelarchiv. Maxima in messaging activity and related real-life events were identified manually. B. Number of messages in the time period between January 1, 2022 and the end of the observation period (blue) and number of messages posted per person per day (green). The numbers in parentheses indicate the position of the events highlighted in panel A.
4.4 Activity distribution
A common aspect of social media platforms is high activity concentration, where few drivers (authors, chats) are responsible for most of the content/attention [17]. The Schwurbelarchiv is no different. Grouping the messages by unique authors and chats (Fig. 1C), we see that the top 10% of chats aggregate 73.3% of all messages and the top 10% authors contribute 94.6% of all messages. This is reflected in the very high Gini indexes of G=0.83 (chats) and G=0.95 (authors). For comparison, Gini coefficients reported for attention metrics on Twitter are in a similar range with G=0.94 for followers and G=0.90 for retweets [18]. It is, however, important to mention that these measure different quantities than posting activity per chat and are therefore not directly comparable. For Telegram and other messaging platforms, comparable activity concentration metrics are, to our knowledge, not available in the published literature.
The probability distributions of messages and unique authors per chat (Fig. 1D) have very wide ranges, with the largest chats having on the order of $10^{4}$ authors and $10^{6}$ messages. Due to the uncertainty around the dataset sampling scheme, we refrain from e.g. fitting power-law distributions as the sampling can heavily bias both measured exponents and model selection of the best-fitted distribution (e.g. power-law vs. lognormal). Regardless, the distributions demonstrate fat or heavy tails.
@@ 4.5 Validation and Completeness of the Dataset
In light of the absence of a structured database of Telegram content that can be used to retrieve a comprehensive dataset for a given query, it is important to assess the completeness of collected Telegram datasets and its relatedness to the targeted scope, e.g. German language conspiracy related discourse. Since Telegram does not provide a searchable index of public chats nor any official record of chats about specific topics, no ground truth exists with which the completeness of any Telegram dataset could be measured. All existing Telegram datasets therefore rely on manual search and selection or snowball sampling. As discussed in Section “Data Collection” above, snowball sampling is systematically biased toward including chats with higher forwarding activity, meaning that smaller or more insular chats are likely underrepresented in the dataset. The metrics we report can therefore provide only an approximate assessment of completeness. To this end, we consider alternative means of studying completeness, using the information that is available about the data set and in the published literature: The anonymous collector of the data set states that the Schwurbelarchiv contains conspiracy-related content. We start by validating this claim via classification of message content. Next, and motivated by the reported frequent message deletions on Telegram [15], we gauge the probability of content deletions to have taken place before data was collected given the data collection approach employed. Next, we attempt to gauge the extent to which the data set covers the discourse that is in scope of the sampling. While no searchable index of existing content exists, forwarded content provides a means of assessing the proportion of the targeted discourse that is included in the sample by analysing the presence of original messages and forwarded messages: if many forwarded messages are not present in the dataset as original messages, this is evidence that a substantial part of the discourse is not captured. Lastly, we employ a complementary approach to assess completeness of the data set by looking at the chats that the Schwurbelarchiv includes. Since there is no official record of all German-speaking Telegram chats dedicated to conspiracy-related content we compare the chats in the dataset with chats in similar datasets, specifically those by Zehring and Domahidi [9], Mohr [13] and Baumgartner et al. [3] introduced earlier.
4.5.1 Conspiracy Content in Telegram Chats
As described in the Methods Section above, the dataset was collected anonymously through snowball sampling of conspiracy-related Telegram groups with the primary aim of capturing conspiracy-related content on Telegram. However, other than manual curation by the data collector, no validation was conducted to ensure that the dataset effectively captured conspiracy content, nor was there an assessment of the proportion of messages within the chats that actually contain such content.
To validate the scope of the dataset, we applied a fine-tuned BERT model developed by Pustet et al. [10], which is designed to detect conspiracy-related content in German Telegram messages. This model was applied to all messages exceeding 5 characters in length. The results showed that 18.4% of chats (1055 chats) do not contain conspiracy-related content. However, these chats account for only 0.5% of all messages longer than 150 characters. This indicates that the majority of the dataset indeed contains conspiracy-related material.
@@ 4.5.2 Deleted messages
Assessing the completeness of the dataset on a message level involves examining the potential for missing or deleted messages, which is closely tied to the data collection approach. There is no published metadata on the time of data collection. In particular, the start of the streaming process (after the initial bulk download upon discovery) for each chat is not known. However, the difference between bulk download and streaming has direct consequences for the expected completeness of messages in the dataset at different points in time: In the bulk download phase, it is probable that the dataset is missing messages that were deleted by authors prior to the download and that the percentage of missing messages increases further back in time from the time of the download. These deleted messages have no placeholders in the dataset and are therefore not visible. In contrast, during the continual streaming phase, messages are captured in near real-time, therefore reducing the probability of missing messages caused by user deletions. We were not able to detect the transition point where the data collection method changed from bulk to stream download for each chat. Therefore, we cannot guarantee the completeness of the dataset on a message level with respect to message deletions, since we cannot indicate the presence or fraction of deleted messages. However, it is probable that message deletions are less likely to have occurred for the later years contained in the dataset and that the Schwurbelarchiv is less affected by message deletions than other comparable data sets.
Even though our data set likely has less deleted messages, researchers using the data should still consider the systematic bias and possibly skewed results in computational analysis introduced by message deletions, as stated by Buehling [15]. These findings indicate a key limitation of our dataset, since the lack of metadata detailing the data collection process prevents us from determining the extent to which deleted messages may have influenced its completeness.
@@ 4.5.3 Forwarded Messages
To further assess the completeness of conspiracy-related chats in the dataset, we performed a comparative analysis of forwarded and original messages. In the context of Telegram, a forwarded message refers to a message which has been forwarded and is marked as such in the chat. These forwarded messages often serve as proxies for content influence, reflecting the flow of information between chats. We calculated the proportion of forwarded messages which are also present in the dataset as original messages. This metric indicates the degree to which the original sources of forwarded messages are represented in the dataset. If this value were 100%, it would imply that all forwarded messages are also present in the original chat the message was posted in, achieving complete coverage of the source population that is in scope of the sampling. In our dataset, this proportion was calculated as the fraction of forwarded messages whose original sender (fwd_author) and original posting time (fwd_posting_date in UTC) jointly match an original (non-forwarded) message in the dataset (matching on author and posting_date after UTC normalisation). In the dataset this value is 54.1%. This finding suggests that we cover about half of the German-language conspiracy related discourse on Telegram in the observation time frame, highlighting the dataset’s limitations in capturing the full extent of the discourse. We interpret this proportion as a rough estimate of coverage. A higher or lower value is not inherently desirable, as it reflects the scope of the sampled forwarding network rather than the quality of the data contained.
4.5.4 Chats Covered
@@
Lastly, we also compare the contents of the Schwurbelarchiv to the chats collected by Zehring and Domahidi [9], Mohr [13] and Baumgartner et al. [3]. The dataset by Zehring and Domahidi [9], consists of 578 chats and covers the time period October 28, 2015, to January 3, 2022. Among these chats, there are two that we have excluded from our dataset due to their classification as dump chats. Out of the remaining chats, 121 are also present in our dataset, representing approximately 22% of Zehring’s chats and 2% of the chats contained in the Schwurbelarchiv. Although this overlap is smaller than expected given the similar selection criteria and seed topics, several factors may account for the discrepancy: Zehring and Domahidi [9] concentrated their seed selection on groups with substantial societal influence, including opinion leaders, politicians, and well-known conspiracy theorists. In contrast, the Schwurbelarchiv dataset includes numerous smaller, more specialized groups often focused on local issues. For example, groups such as Corona Rebellen have multiple sub-groups tailored to different cities and regions, reflecting a tendency towards decentralization in the Schwurbelarchiv dataset.
@@
The subset of German-language chats in the dataset collected by Mohr [13] which was used for comparison here consists of 22,968 chats. Among these, 46 have been categorised by us as dump chats and have been excluded. Another 801 of the chats in this dataset are also contained in the Schwurbelarchiv. This corresponds to 13% of the chats we have in our dataset, while only comprising 3% of chats contained in the dataset collected by Mohr [13].
@@
The Pushshift dataset [3] is the most comprehensive public dataset to date, encompassing over 27,000 chats from various countries. Of these, 78 chats (1.3%) overlap with the Schwurbelarchiv. None of the excluded dump chats in the Schwurbelarchiv are present in the Pushshift dataset. This low overlap is likely due to linguistic and regional differences in the selection criteria: while the Schwurbelarchiv and datasets by Zehring and Domahidi [9] and Mohr [13] focus on German-speaking chats, the Pushshift dataset has a more global scope. In addition, the temporal scope of data collection between the Pushshift data set (2015-2019) and the Schwurbelarchiv (2016-2022) differs substantially.
Table 3 summarizes the number of chats in each dataset and their overlap with the Schwurbelarchiv, reported both in absolute numbers and as a percentage of Schwurbelarchiv chats. Furthermore, the overlap between the Zehring and Domahidi [9] and Mohr [13] datasets is 369 chats, corresponding to 66% and 1.6% of the content of each of these datasets, respectively. A total of 102 chats are present in all three data sets. Additionally, 39 chats are shared between the Mohr [13] and Pushshift datasets, while no chats overlap between the datasets by Zehring and Domahidi [9] and Pushshift.
Dataset Description Number of Chats Overlap Overlap in %
Mohr [13] COVID-19 misinformation; multilingual; keyword-guided snowball sampling 22,968 801 13.0%
Zehring & Domahidi [9] German Querdenken movement; far-right and conspiracy content 578 121 2.0%
Pushshift [3] General; multilingual; seeded from right-wing extremism and cryptocurrency channels 27,801 78 1.3%
@@
Additionally, we compared our dataset with the channels and groups monitored by the Machine Against the Rage initiative by BAG Gegen Hass im Netz [19], which shows network patterns of forwarded messages within the ideological spectrum of conspiracy theories and extremism. We found an overlap of 212 chats with the Schwurbelarchiv. The initiative’s monitoring begins right after our observation period in September 2022 and therefore allows for contextualization of these groups’ further embeddedness in the German speaking Telegram network.